a body is moving with uniform velocity of 8 m/s when the body just crosses another body the secound one satrts and mooves with unform acceleratan of 4 meter per sec sqauare the time after witch to body will meet is?

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
manas
5 years ago
here, 1st body is moving with uniform velocity of 8m/s.
W.K.T, 
velocity=displacement/time
8=s/t
so,
s=8t
 
for second body,
it starts after 1st body crosses it,
u=o
a=4m/s^2
so,
s=1/2.4.t^2
since displacement of both bodies are same
8t=1/2.4.t^2
by solving we get,
8t=2t^2
so t=4secs